OpenStudy (anonymous):

Solve 2(x - 4) - 6(2x - 3) = 0 A. x = -4 B. x = -1 C. x = 0 D. x = 1

OpenStudy (anonymous):

2x - 8 - 12x + 18 = 0 -10x = -18 + 8 -10x = -10 x = 1 So option D.